Need help with this question

A particle of mass m is in the one-dimensional potential a. Argue, based on dimensional arguments dhat the enersy levelust be of the form With α, β, γ being numerical coefficients and c(n) being a positive numerical coefficient that depends on the energy levels. Determine the values of a, β and γ.
Answer #1

E_{n} has dimension of energy as it represents n-th energy level. So dimension of E_{n} is [MIT-2].

As E_{n} = c(n) hslash^ {alpha} m^{eta}F^{gamma} and c(n) is a numerical coefficient,

So, dimension of hslash^ {alpha} m^{eta}F^{gamma} is same as dimension of E_{n}.

We know dimension of hslash , m , Fare [MIT-] , [M] and [MLT^{-2}] respectively.

So, dimension of hslash^ {alpha} m^{eta}F^{gamma} = [ML^2T^{-1}]^{alpha}[M]^{eta}[MLT^{-1}]^{gamma} = dimension of E_{n}.

Rightarrow [M^{alpha + eta + gamma}L^{2alpha +gamma}T^{-alpha-2gamma}] = [MIT-2]

Comparison of powers on both side leads to-

alpha +eta +gamma =1

2a+1=2

-alpha -2gamma = -2

Solving above three equations we can have,

alpha =2/3, eta= -1/3, gamma= 2/3

Hope this helps.

Good Luck.